### v2

@@ 2.1 The mutual information poset and Klein’s condition
Having defined KC, we now review how it can be conveniently rephrased in terms of a partially ordered structure. Recall that a poset $\mathcal{P}=(X,\preceq)$ is defined as a set $X$ with a binary relation $\preceq$ that is reflexive, antisymmetric, and transitive.181818 That is, for all $x,y,z\in X$ we have (i) $x\preceq x$; (ii) if $x\preceq y$ and $y\preceq x$, then $x=y$; and finally (iii) if $x\preceq y$ and $y\preceq z$, then $x\preceq z$. We define the mutual information poset (MI-poset) as follows.
@@
It then follows from the definition of the partial order in the MI-poset (15) and the definition of KC given in (13) that a face $\mathscr{F}$ of the SACN is a KC-face only if $\mathcal{V}(\mathscr{F})$ is a down-set. It is important to notice that given an arbitrary down-set $\mathcal{D}$ of the MI-poset, it is not necessarily the case that there exists a face $\mathscr{F}$ such that $\mathcal{V}(\mathscr{F})=\mathcal{D}$. An analogous statement also holds for arbitrary subsets $\mathcal{X}\subseteq\mathcal{E}$, and the underlying reason is in fact the same: Such arbitrary $\mathcal{D}$ and $\mathcal{X}$ do not necessarily respect the dependences among SA instances. For now, however, it is sufficient to notice that by focusing on down-sets of the MI-poset, we can restrict the search space for the brute force computation of $\mathscr{R}_{{}_{\text{SSA}}}^{{}^{{\sf N}}}$ from the power-set $2^{\mathcal{E}}$ to the lattice of down-sets191919 The set of all down-sets of any poset, with partial order given by inclusion, has the structure of a lattice, where meet and join correspond respectively to intersection and union birkhoff1967lattice . of the MI-poset, which is already significantly smaller. For example, for ${\sf N}=3$ we have $|2^{\mathcal{E}}|=2^{25}=33,\!554,\!432$, while the down-set lattice has cardinality $11,\!422$.
@@ 2.2 The Bell pairs theorem and additional constraints
where $\underline{\mathscr{J}}^{c}\equiv[\![{\sf N}]\!]\setminus\underline{\mathscr{J}}$, and we utilized the assumption that the system on all $[\![{\sf N}]\!]$ parties is pure.202020 This implies that ${\sf S}_{[\![{\sf N}]\!]}={\sf S}_{\varnothing}=0$ and ${\sf S}_{\underline{\mathscr{J}}^{c}}={\sf S}_{\underline{\mathscr{J}}}$. Consider now a face $\mathscr{F}$ of the SACN such that $\mathcal{V}(\mathscr{F})$ includes one such element. Assuming for simplicity that $\underline{\mathscr{J}}$ does not include the purifier $\ell=0$ (if it does we can simply repeat the same argument for $\underline{\mathscr{J}}^{c}$), we can then denote $\underline{\mathscr{J}}$ by just $\mathscr{J}$ and have ${\sf S}_{\mathscr{J}}=0$ for any vector $\vec{{\sf S}}\in\mathscr{F}$. If $\vec{{\sf S}}$ is realizable by some density matrix $\rho_{{\sf N}}$, then we can write it as
@@
where |$\ket{0}\bra{0}_{\mathscr{J}}$ is a “completely decorrelated” state for subsystem $\mathscr{J}$212121 In other words, |$\ket{0}_{\mathscr{J}}$ is a tensor product of states for the individual parties in $\mathscr{J}$. If $\mathscr{F}$ is an extreme ray, any entropy vector $\vec{\sf S}(\rho_{\sf N})$ generating it cannot be the sum of two distinct nontrivial vectors within the SAC. Therefore, the only possibility is that the first term is the null vector, which implies
@@ 3 The algorithm
In this section, we present the algorithm that we will use to find $\mathscr{R}_{{}_{\text{KC}}}^{6}$. Since nothing about this algorithm is specific to the set-up of the subadditivity cone described in the previous section, we will present it more generally, such that it can be applied to other contexts without any adaptation. In general terms, our algorithm can be used when one has a polyhedral cone specified by a set of linear inequalities with partial order among them, and one wants to find only a particular subset of extreme rays, characterized by the fact that for each extreme ray the set of inequalities which are saturated is a down-set in the poset.222222 To make the algorithm more robust, we will make a few additional assumptions, which will be presented in detail below.
We anticipate that this algorithm is typically most useful in situations where the poset has “large height and small width.” In the limiting situation where the poset is a chain, our algorithm is very efficient (although somewhat trivial). In the opposite situation, where the poset is an antichain, it is instead very inefficient since it reduces to the brute force strategy mentioned in the Introduction, and one should instead use one of the standard algorithms for “polyhedral representation conversion” (although in general even these algorithms are not efficient) survey1980 .232323 Notice however that this inefficiency is related to the fact that the subset of extreme rays found by the algorithm becomes the full set of all extreme rays of the cone. In intermediate situations, the efficiency will depend on the details of the problem at hand. Nevertheless, even for the specific problem of finding $\mathscr{R}_{{}_{\text{KC}}}$ indicated in the preceding section, where the height (approximately ${\sf N}$) is exponentially smaller by a factor of roughly $2^{D}/\sqrt{N}$ than the width (approximately $\binom{{\sf N}}{{\sf N}/2}$), the algorithm remains remarkably useful. In any case, a feature of our algorithm is the flexibility of how it can be used in practice, and the possibility of easily combining it with another conversion algorithm. Essentially, one first runs our algorithm to extract a collection of lower-dimensional faces of the cone that is guaranteed to contain all the “down-set extreme rays.” Then one can find all the extreme rays for each of these faces using standard techniques and finally check which ones correspond to down-sets (which is a straightforward computation). Finally, we stress that even in this situation, the algorithm will provide useful information about “excluded regions” on the individual faces that in many cases can be used to simplify the derivation of the down-set extreme rays using standard conversion algorithms.
@@ 3.1 Set-up
Consider a vector space $\mathbb{R}^{n}$, along with a finite collection of linear inequalities that specify a full-dimensional pointed cone $\mathcal{C}$242424 Any cone specified by a finite set of inequalities is necessarily convex and polyhedral. If the cone is not full-dimensional we imagine to first “dimensionally reduce” the set-up to the subspace spanned by the cone. If it is not pointed, it is the Minkowski sum of a lower-dimensional cone and a linear subspace, and we can focus on this lower-dimensional cone. We will write an inequality as ${\sf{E}}_{i}(\vec{v})\geq 0$, where ${\sf{E}}_{i}$ is an element of the dual space of $\mathbb{R}^{n}$ and $\vec{v}$ denotes an arbitrary vector in $\mathbb{R}^{n}$. Let $\mathcal{E}$ denote the set of dual vectors ${\sf{E}}_{i}$, with $i\in[k]$ and $k=|\mathcal{E}|$. A face $\mathscr{F}$ of $\mathcal{C}$ is defined as the intersection of $\mathcal{C}$ with a hyperplane $\mathbb{H}$ in $\mathbb{R}^{n}$ such that $\mathcal{C}$ is entirely contained in one of the two (closed) half-spaces specified by $\mathbb{H}$252525 The full cone is also a face, obtained in the degenerate case where ${\sf{E}}_{i}=\vec{0}$. In what follows we will ignore this face and always assume ${\sf{E}}_{i}\neq\vec{0}$. A codimension-1 face is called a facet, and a $1$-dimensional face an extreme ray.
@@
In general, given an arbitrary collection of inequalities that specifies a full-dimensional pointed cone, it is not always the case that each inequality determines a facet. An inequality ${\sf{E}}_{i}(\vec{v})\geq 0$ such that $\mathscr{F}_{i}$ is not a facet is called a redundant inequality. Geometrically, a redundant inequality ${\sf{E}}_{i}(\vec{v})\geq 0$ is characterized by the fact that for any $\vec{v}_{*}\in\mathcal{C}$ such that ${\sf{E}}_{i}(\vec{v}_{*})=0$, we necessarily have ${\sf{E}}_{j}(\vec{v}_{*})=0$ for some other ${\sf{E}}_{j}$ in $\mathcal{E}$. Algebraically, it is characterized by the fact that it is a positive linear combination of other inequalities in the set. Thus, if ${\sf{E}}_{i}(\vec{v})\geq 0$ is redundant, then the system of inequalities obtained by “ignoring” ${\sf{E}}_{i}$ (i.e., the set of inequalities specified by =$\mathcal{E}^{\prime}=\mathcal{E}\setminus\{{\sf{E}}_{i}\}$) carves out the same cone $\mathcal{C}$. Given an arbitrary set of inequalities, one can use a linear program to check if a specific inequality is redundant. Indeed, there are efficient algorithms which reduce the set of inequalities to a minimum subset that specifies the same cone, where each inequality corresponds to a facet Fukuda15 . We therefore assume that each ${\sf{E}}_{i}\in\mathcal{E}$ corresponds to a non-redundant inequality.262626 This assumption is mainly made here for the purpose of defining the cone and its symmetries. However, we will introduce certain redundant inequalities in §3.3 to specify a region of the cone that we want to ignore. This is why at the end of the previous section we chose to keep the maximal elements of the MI-poset.
@@
This is the subset of $\mathcal{E}$ whose elements specify the inequalities which are saturated by the vectors in the interior of $\mathscr{F}$ (notice that this set is independent from the choice of vector $\vec{v}\in\text{int}(\mathscr{F})$). If $\mathcal{V}(\mathscr{F})$ is a down-set in $\mathcal{P}$, we will say that $\mathscr{F}$ is a down-set face (D-face). Our goal will be to find the 1-dimensional D-faces, or down-set extreme rays (D-ERs), of $\mathcal{C}$272727 For the case where $\mathcal{P}$ is the MI-poset, D-faces and D-ERs are precisely the KC-faces and KC-ERs, respectively.
@@
Since the geometric and order structures of our set-up are independent from each other, an element of $G_{\mathcal{C}}$ in general might map a D-face to another face that is not a D-face. We assume now that the “structure” of the down-set corresponding to a D-face is important for the problem we want to solve with our algorithm,313131 Otherwise, there is no need to introduce the poset structure of $\mathcal{E}$ in the first place. so we want to consider the group of transformations that preserve this structure. To make this precise, we specify in what sense two down-sets can be considered structurally equivalent. The set of down-sets of $\mathcal{P}$, ordered by inclusion, forms a lattice $\mathcal{L}_{\mathcal{P}}$, which we shall call the down-set lattice birkhoff1967lattice . As for the case of $\mathcal{L}_{\mathcal{C}}$, we then define the combinatorial automorphism group of the poset $G_{\mathcal{P}}$ as the largest subgroup of $\text{Sym}(k)$ that preserves this lattice. Two down-sets of $\mathcal{P}$ will be considered equivalent if they are related by such a permutation.
In general, $\mathcal{L}_{\mathcal{C}}$ and $\mathcal{L}_{\mathcal{P}}$, and therefore $G_{\mathcal{C}}$ and $G_{\mathcal{P}}$, are unrelated. Since we want to preserve both lattices, we are then interested in the largest common subgroup323232 The set $\mathcal{L}_{\mathcal{C}\mathcal{P}}=\mathcal{L}_{\mathcal{C}}\cap\mathcal{L}_{\mathcal{P}}$ also admits a lattice structure, since it is the set of closed subsets of $\mathcal{E}$ under the closure operator $\text{cl}_{\text{FD}}$ (defined in the next section). However, in general $\mathcal{L}_{\mathcal{C}\mathcal{P}}$ is not a sublattice of either $\mathcal{L}_{\mathcal{C}}$ or $\mathcal{L}_{\mathcal{P}}$, though any $g\in G_{\mathcal{C}\mathcal{P}}$ is still an automorphism of $\mathcal{L}_{\mathcal{C}\mathcal{P}}$ for the same reason explained in Footnote 30.
@@
Finally, we also consider the possibility that in the problem we want to solve with our algorithm, we are only interested in finding the extreme rays that do not belong to a specified collection $\widehat{\mathcal{C}}$ of faces of $\mathcal{C}$. We will call $\widehat{\mathcal{C}}$ the excluded region of $\mathcal{C}$ and assume that it is closed under the action of $G$. Specifically, each face $\mathscr{F}\in\widehat{\mathcal{C}}$ can be described by its corresponding set of saturated inequalities $\mathcal{V}(\mathscr{F})$, and using $\mathcal{V}(\mathscr{F})$ it is straightforward to close its orbit under the action of $G$333333 Notice that we are not making any assumption about the specific faces that appear in $\widehat{\mathcal{C}}$, and in general some of them can be facets. In this case it might be possible to simplify the problem by ignoring some of the inequalities to obtain a larger cone that is guaranteed to include all extreme rays of $\mathcal{C}$ which are not in $\widehat{\mathcal{C}}$. We will further comment on this possibility in the next section, since in one variation of our algorithm we will apply this strategy to lower-dimensional faces generated during the computation. While for a given specific problem the excluded region can be empty (no face is excluded), this notion, appropriately generalized to lower-dimensional faces, will play a key role during the computation. Since the algorithm proceeds by iteration, exploring at each step orbits of faces of various dimensionalities, we will use the excluded region to keep track of faces that have already been explored, or that are guaranteed to not include any new D-ER (more on this below), thereby allowing us to ignore them at subsequent steps of the computation.

### v3

@@ 2.1 The mutual information poset and Klein’s condition
Having defined KC, we now review how it can be conveniently rephrased in terms of a partially ordered structure. Recall that a poset $\mathcal{P}=(X,\preceq)$ is defined as a set $X$ with a binary relation $\preceq$ that is reflexive, antisymmetric, and transitive.171717 That is, for all $x,y,z\in X$ we have (i) $x\preceq x$; (ii) if $x\preceq y$ and $y\preceq x$, then $x=y$; and finally (iii) if $x\preceq y$ and $y\preceq z$, then $x\preceq z$. We define the mutual information poset (MI-poset) as follows.
@@
It then follows from the definition of the partial order in the MI-poset (15) and the definition of KC given in (13) that a face $\mathscr{F}$ of the SACN is a KC-face only if $\mathcal{V}(\mathscr{F})$ is a down-set. It is important to notice that given an arbitrary down-set $\mathcal{D}$ of the MI-poset, it is not necessarily the case that there exists a face $\mathscr{F}$ such that $\mathcal{V}(\mathscr{F})=\mathcal{D}$. An analogous statement also holds for arbitrary subsets $\mathcal{X}\subseteq\mathcal{E}$, and the underlying reason is in fact the same: Such arbitrary $\mathcal{D}$ and $\mathcal{X}$ do not necessarily respect the dependences among SA instances. For now, however, it is sufficient to notice that by focusing on down-sets of the MI-poset, we can restrict the search space for the brute force computation of $\mathscr{R}_{{}_{\text{SSA}}}^{{}^{{\sf N}}}$ from the power-set $2^{\mathcal{E}}$ to the lattice of down-sets181818 The set of all down-sets of any poset, with partial order given by inclusion, has the structure of a lattice, where meet and join correspond respectively to intersection and union birkhoff1967lattice . of the MI-poset, which is already significantly smaller. For example, for ${\sf N}=3$ we have $|2^{\mathcal{E}}|=2^{25}=33,\!554,\!432$, while the down-set lattice has cardinality $11,\!422$.
@@ 2.2 The Bell pairs theorem and additional constraints
where $\underline{\mathscr{J}}^{c}\equiv[\![{\sf N}]\!]\setminus\underline{\mathscr{J}}$, and we utilized the assumption that the system on all $[\![{\sf N}]\!]$ parties is pure.191919 This implies that ${\sf S}_{[\![{\sf N}]\!]}={\sf S}_{\varnothing}=0$ and ${\sf S}_{\underline{\mathscr{J}}^{c}}={\sf S}_{\underline{\mathscr{J}}}$. Consider now a face $\mathscr{F}$ of the SACN such that $\mathcal{V}(\mathscr{F})$ includes one such element. Assuming for simplicity that $\underline{\mathscr{J}}$ does not include the purifier $\ell=0$ (if it does we can simply repeat the same argument for $\underline{\mathscr{J}}^{c}$), we can then denote $\underline{\mathscr{J}}$ by just $\mathscr{J}$ and have ${\sf S}_{\mathscr{J}}=0$ for any vector $\vec{{\sf S}}\in\mathscr{F}$. If $\vec{{\sf S}}$ is realizable by some density matrix $\rho_{{\sf N}}$, then we can write it as
@@
where |$\ket{0}\bra{0}_{\mathscr{J}}$ is a “completely decorrelated” state for subsystem $\mathscr{J}$202020 In other words, |$\ket{0}_{\mathscr{J}}$ is a tensor product of states for the individual parties in $\mathscr{J}$. If $\mathscr{F}$ is an extreme ray, any entropy vector $\vec{\sf S}(\rho_{\sf N})$ generating it cannot be the sum of two distinct nontrivial vectors within the SAC. Therefore, the only possibility is that the first term is the null vector, which implies
@@ 3 The algorithm
In this section, we present the algorithm that we will use to find $\mathscr{R}_{{}_{\text{KC}}}^{6}$. Since nothing about this algorithm is specific to the set-up of the subadditivity cone described in the previous section, we will present it more generally, such that it can be applied to other contexts without any adaptation. In general terms, our algorithm can be used when one has a polyhedral cone specified by a set of linear inequalities with partial order among them, and one wants to find only a particular subset of extreme rays, characterized by the fact that for each extreme ray the set of inequalities that are saturated is a down-set in the poset.212121 To make the algorithm more robust, we will make a few additional assumptions, which will be presented in detail below.
We anticipate that this algorithm is typically most useful in situations where the poset has “large height and small width.” In the limiting situation where the poset is a chain, our algorithm is very efficient (although somewhat trivial). In the opposite situation, where the poset is an antichain, it is instead very inefficient since it reduces to the brute force strategy mentioned in the Introduction, and one should instead use one of the standard algorithms for “polyhedral representation conversion” (although in general even these algorithms are not efficient) survey1980 .222222 Notice however that this inefficiency is related to the fact that the subset of extreme rays found by the algorithm becomes the full set of all extreme rays of the cone. In intermediate situations, the efficiency will depend on the details of the problem at hand. Nevertheless, even for the specific problem of finding $\mathscr{R}_{{}_{\text{KC}}}$ indicated in the preceding section, where the height (approximately ${\sf N}$) is exponentially smaller by a factor of roughly $2^{D}/\sqrt{N}$ than the width (approximately $\binom{{\sf N}}{{\sf N}/2}$), the algorithm remains remarkably useful. In any case, a feature of our algorithm is the flexibility of how it can be used in practice, and the possibility of easily combining it with another conversion algorithm. Essentially, one first runs our algorithm to extract a collection of lower-dimensional faces of the cone that is guaranteed to contain all the “down-set extreme rays.” Then one can find all the extreme rays for each of these faces using standard techniques and finally check which ones correspond to down-sets (which is a straightforward computation). Finally, we stress that even in this situation, the algorithm will provide useful information about “excluded regions” on the individual faces that in many cases can be used to simplify the derivation of the down-set extreme rays using standard conversion algorithms.
@@ 3.1 Set-up
Consider a vector space $\mathbb{R}^{n}$, along with a finite collection of linear inequalities that specify a full-dimensional pointed cone $\mathcal{C}$232323 Any cone specified by a finite set of inequalities is necessarily convex and polyhedral. If the cone is not full-dimensional we imagine to first “dimensionally reduce” the set-up to the subspace spanned by the cone. If it is not pointed, it is the Minkowski sum of a lower-dimensional cone and a linear subspace, and we can focus on this lower-dimensional cone. We will write an inequality as ${\sf{E}}_{i}(\vec{v})\geq 0$, where ${\sf{E}}_{i}$ is an element of the dual space of $\mathbb{R}^{n}$ and $\vec{v}$ denotes an arbitrary vector in $\mathbb{R}^{n}$. Let $\mathcal{E}$ denote the set of dual vectors ${\sf{E}}_{i}$, with $i\in[k]$ and $k=|\mathcal{E}|$. A face $\mathscr{F}$ of $\mathcal{C}$ is defined as the intersection of $\mathcal{C}$ with a hyperplane $\mathbb{H}$ in $\mathbb{R}^{n}$ such that $\mathcal{C}$ is entirely contained in one of the two (closed) half-spaces specified by $\mathbb{H}$242424 The full cone is also a face, obtained in the degenerate case where ${\sf{E}}_{i}=\vec{0}$. In what follows we will ignore this face and always assume ${\sf{E}}_{i}\neq\vec{0}$. A codimension-1 face is called a facet, and a $1$-dimensional face an extreme ray.
@@
In general, given an arbitrary collection of inequalities that specifies a full-dimensional pointed cone, it is not always the case that each inequality determines a facet. An inequality ${\sf{E}}_{i}(\vec{v})\geq 0$ such that $\mathscr{F}_{i}$ is not a facet is called a redundant inequality. Geometrically, a redundant inequality ${\sf{E}}_{i}(\vec{v})\geq 0$ is characterized by the fact that for any $\vec{v}_{*}\in\mathcal{C}$ such that ${\sf{E}}_{i}(\vec{v}_{*})=0$, we necessarily have ${\sf{E}}_{j}(\vec{v}_{*})=0$ for some other ${\sf{E}}_{j}$ in $\mathcal{E}$. Algebraically, it is characterized by the fact that it is a positive linear combination of other inequalities in the set. Thus, if ${\sf{E}}_{i}(\vec{v})\geq 0$ is redundant, then the system of inequalities obtained by “ignoring” ${\sf{E}}_{i}$ (i.e., the set of inequalities specified by =$\mathcal{E}^{\prime}=\mathcal{E}\setminus\{{\sf{E}}_{i}\}$) carves out the same cone $\mathcal{C}$. Given an arbitrary set of inequalities, one can use a linear program to check if a specific inequality is redundant. Indeed, there are efficient algorithms which reduce the set of inequalities to a minimum subset that specifies the same cone, where each inequality corresponds to a facet Fukuda15 . We therefore assume that each ${\sf{E}}_{i}\in\mathcal{E}$ corresponds to a non-redundant inequality.252525 This assumption is mainly made here for the purpose of defining the cone and its symmetries. However, we will introduce certain redundant inequalities in §3.3 to specify a region of the cone that we want to ignore. This is why at the end of the previous section we chose to keep the maximal elements of the MI-poset.
@@
This is the subset of $\mathcal{E}$ whose elements specify the inequalities which are saturated by the vectors in the interior of $\mathscr{F}$ (notice that this set is independent from the choice of vector $\vec{v}\in\text{int}(\mathscr{F})$). If $\mathcal{V}(\mathscr{F})$ is a down-set in $\mathcal{P}$, we will say that $\mathscr{F}$ is a down-set face (D-face). Our goal will be to find the 1-dimensional D-faces, or down-set extreme rays (D-ERs), of $\mathcal{C}$262626 For the case where $\mathcal{P}$ is the MI-poset, D-faces and D-ERs are precisely the KC-faces and KC-ERs, respectively.
@@
Since the geometric and order structures of our set-up are independent from each other, an element of $G_{\mathcal{C}}$ in general might map a D-face to another face that is not a D-face. We assume now that the “structure” of the down-set corresponding to a D-face is important for the problem we want to solve with our algorithm,303030 Otherwise, there is no need to introduce the poset structure of $\mathcal{E}$ in the first place. so we want to consider the group of transformations that preserve this structure. To make this precise, we specify in what sense two down-sets can be considered structurally equivalent. The set of down-sets of $\mathcal{P}$, ordered by inclusion, forms a lattice $\mathcal{L}_{\mathcal{P}}$, which we shall call the down-set lattice birkhoff1967lattice . As for the case of $\mathcal{L}_{\mathcal{C}}$, we then define the combinatorial automorphism group of the poset $G_{\mathcal{P}}$ as the largest subgroup of $\text{Sym}(k)$ that preserves this lattice. Two down-sets of $\mathcal{P}$ will be considered equivalent if they are related by such a permutation.
In general, $\mathcal{L}_{\mathcal{C}}$ and $\mathcal{L}_{\mathcal{P}}$, and therefore $G_{\mathcal{C}}$ and $G_{\mathcal{P}}$, are unrelated. Since we want to preserve both lattices, we are then interested in the largest common subgroup313131 The set $\mathcal{L}_{\mathcal{C}\mathcal{P}}=\mathcal{L}_{\mathcal{C}}\cap\mathcal{L}_{\mathcal{P}}$ also admits a lattice structure, since it is the set of closed subsets of $\mathcal{E}$ under the closure operator $\text{cl}_{\text{FD}}$ (defined in the next section). However, in general $\mathcal{L}_{\mathcal{C}\mathcal{P}}$ is not a sublattice of either $\mathcal{L}_{\mathcal{C}}$ or $\mathcal{L}_{\mathcal{P}}$, though any $g\in G_{\mathcal{C}\mathcal{P}}$ is still an automorphism of $\mathcal{L}_{\mathcal{C}\mathcal{P}}$ for the same reason explained in Footnote 29.
@@
Finally, we also consider the possibility that in the problem we want to solve with our algorithm, we are only interested in finding the extreme rays that do not belong to a specified collection $\widehat{\mathcal{C}}$ of faces of $\mathcal{C}$. We will call $\widehat{\mathcal{C}}$ the excluded region of $\mathcal{C}$ and assume that it is closed under the action of $G$. Specifically, each face $\mathscr{F}\in\widehat{\mathcal{C}}$ can be described by its corresponding set of saturated inequalities $\mathcal{V}(\mathscr{F})$, and using $\mathcal{V}(\mathscr{F})$ it is straightforward to close its orbit under the action of $G$323232 Notice that we are not making any assumption about the specific faces that appear in $\widehat{\mathcal{C}}$, and in general some of them can be facets. In this case it might be possible to simplify the problem by ignoring some of the inequalities to obtain a larger cone that is guaranteed to include all extreme rays of $\mathcal{C}$ which are not in $\widehat{\mathcal{C}}$. We will further comment on this possibility in the next section, since in one variation of our algorithm we will apply this strategy to lower-dimensional faces generated during the computation. While for a given specific problem the excluded region can be empty (no face is excluded), this notion, appropriately generalized to lower-dimensional faces, will play a key role during the computation. Since the algorithm proceeds by iteration, exploring at each step orbits of faces of various dimensionalities, we will use the excluded region to keep track of faces that have already been explored, or that are guaranteed to not include any new D-ER (more on this below), thereby allowing us to ignore them at subsequent steps of the computation.